### v4

# Formal conjugacy and asymptotic differential algebra

Abstract
We study conjugacy of formal derivations on fields of generalised power series in characteristic 0 and dimension $1$. Casting the problem of Poincaré resonance in terms of asymptotic differential algebra, we give conditions for conjugacy of parabolic flat $\log$-$\exp$ transseries, flat grid-based transseries, logarithmic transseries, power series with exponents and coefficients in an ordered field, and formal Puiseux series.
Introduction
Conjugating and normalising local analytic diffeomorphisms around fixed points is a classical method for classifying dynamical systems. Its formal version, where the conjugating element is a possibly divergent formal power series, is usually easier to tackle, as it is devoid of convergence issues (see [17]). Yet normalising formal power series may be difficult because of the phenomenon of resonance, as first studied by Poincaré [33] and Dulac [13]. Resonance introduces non-convexity into the conjugacy problem: given three series $f$, $g$ and $h$ where $h$ is closer to $f$ than $g$ from a valuative standpoint, it may be that $f$ and $g$ are conjugate whereas $f$ and $h$ are not (see Section 3.3).
Our main motivation for this paper comes from our interest in first-order properties of certain valued groups [8, 6]. From the model theoretic standpoint, the existence of resonance means that the geometry of definable sets in those structures is too complicated to study. Thus finding contexts in which conjugacy is resonance-free is crucial. We hope to convince the reader that as far as formal normalisation of local objects is concerned, the three following notions, belonging to seemingly disparate domains, are strongly connected:
    non-resonance, as a linear algebraic condition for linearisation of vector fields [17],
    convexity of conjugacy, as a case of tameness of definable sets in valued groups [6],
    asymptotic integration, as a closure property for valued differential fields [34].
In connection [20, 31] with Dulac’s problem (see [14, 25, 15]), the dynamics of Poincaré first-return maps, and the analysis of limit cycles of vector fields via Dulac series, there has been interest recently [32, 30, 29] in normalising formal series which may involve formal exponentials and logarithms of the infinite variable $x$. This is for instance notable in Écalle’s method of linearisation by compensators [17]. For these more general questions, a natural domain of investigation is the field of logarithmic-exponential transseries [11, 16, 12]. It is known how to normalise purely logarithmic transseries [21] of the form $\lambda x+o(x)$, in the hyperbolic case [32], i.e. when $\lambda\neq 1$, and in the parabolic case [30], when $\lambda=1$. There is ongoing work on the hyperbolic and parabolic cases in the more general setting of H-fields [1, 2] equipped with composition laws (see [8, Section 4.1]).
In this paper, we focus on the conjugacy problem for parabolic series in differential valued fields [34]. These include for instance formal Laurent or Puiseux series, $\log$-$\exp$ transseries or grid-based transseries [24], logarithmic transseries [21], or $\omega$-series [5], and complexifications thereof. This choice is not fortuitous but motivated by the connections between the setting of (ordered) asymptotic differential algebra and that of (ordered) valued groups (see [6, Remark 7.27]). One of the difficulties of solving conjugacy equations for formal series endowed with a composition law $\circ$ and a derivation $\partial$ is that this requires a good understanding of the interaction of the composition law with the valuation. On this path, one is confronted with intricate and computationally heavy problems involving Taylor expansions of arbitrarily high orders, monotonicity of the composition law, and mean value inequalities. It is not the least of hindrances that such properties of $\circ$ and $\partial$ may not have been established for the given algebra of formal series.
We circumvent these issues by leveraging the Lie-type correspondence, given by a formal exponential map $\exp$, between near-identity substitutions $f\mapsto f\circ(x+\delta)$ and contracting derivations $f\mapsto g\partial(f)$ on algebras of formal series. We showed [9] that a fraction of the theory of Lie groups applies to such algebras. We focus on contracting derivations on fields of generalised series, i.e. derivations which piecewise increase the valuation of non-trivial elements. On the side of vector fields, contractiveness can be related to nilpotency (see [28]).
Given a direct limit $\mathbb{S}$ of fields of Hahn series with its natural valuation $v$ and a derivation $\partial:f\mapsto f^{\prime}$ on $\mathbb{S}$ that is compatible with the structure of direct limit of fields of series (see Definition 1.1), we consider a group $(\operatorname{Cont}(\partial),\ast)$ introduced in [6] of contracting derivations on $\mathbb{S}$. The group law $\ast$ is a formal Baker-Campbell-Hausdorff operation (see [27, 26, 35, 10]). This group also has a structure of Lie algebra and can be seen as a linearisation of its “Lie group” $\exp(\operatorname{Cont}(\partial))$. The latter is a group, under functional composition, of substitutions. In $(\operatorname{Cont}(\partial),\ast)$, finding approximate solutions of conjugacy equations reduces to finding approximate solutions of linear differential equations of order $1$ (see Lemma 2.1). Using spherical completeness arguments, one can obtain exact solutions by transfinitely iterating the approximation method (see Lemma 2.2). This composition-free framework allows us to easily understand obstructions to conjugacy, and to cast resonance merely as a property of asymptotic differential algebra, i.e. as a property of the valued differential field $(\mathbb{S},v,\partial)$. This gives a simple connection between features of the asymptotic couple [4, Section 9.1] of $\mathbb{S}$, in particular asymptotic integration [4, p 327], and the existence of resonance for conjugacy equations and normal forms. Say that $\mathbb{S}$ has regular asymptotic integration if asymptotic integration on $\mathbb{S}$ is compatible with its structure of direct limit (see Definition 1.2). We prove:
Theorem 1
[Theorem 2.1] Suppose that $\mathbb{S}$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$. Then $f$ and $g$ are conjugate in $\operatorname{Cont}(\partial)$ if and only if v(fg)>v(g)v(f-g)>v(g) and $v(f-g)>v\left(\left(h/gh^{\prime}\right)^{\prime}\right)$ for all $h\in\mathbb{S}^{\times}$ with $v(h)>0$.
Another benefit of our method is that it is independent of the field of scalars $C$, in that the results are preserved under extensions of scalars (see Remark 1.2). Using the Lie-type correspondence, we obtain a more classical reformulation of the conjugacy problem:
Theorem 2
[Theorem 2.2] Suppose that $\mathbb{S}$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ such that v(fg)>v(f)v(f-g)>v(f) and and $v(f-g)>v\left(\left(h/gh^{\prime}\right)^{\prime}\right)$ for all $h\in\mathbb{S}^{\times}$ with $v(h)>0$. Then the derivations $f\partial:=h\mapsto f\partial(h)$ and $g\partial:=h\mapsto g\partial(h)$ are conjugate over $\operatorname{Aut}(\mathbb{S})$, i.e. there is a $\sigma=\exp(h\partial)\in\operatorname{Aut}(\mathbb{S})$ such that $\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}=f\partial$.
In certain cases, the group of automorphisms $\exp(\operatorname{Cont}(\partial))$ is isomorphic to a well-identified group of series under composition. Let $\mathbb{S}$ be the field of transseries whose transmonomials $\mathfrak{m}$ satisfy $v(\mathfrak{m}^{\prime}/\mathfrak{m})+v(x)\geqslant 0$. We identify (Proposition 3.1) the group $\exp(\operatorname{Cont}(\partial))$ for all direct limits of subsystems (Definition 3.1) of the direct system which defines $\mathbb{S}$. Combining this with Theorem 1, we obtain:
Theorem 3
[Theorem 3.2] For all $\delta,\varepsilon\in\mathbb{S}$ with $v(\delta),v(\varepsilon)>v(x)$, the series $x+\delta$ and $x+\varepsilon$ are conjugate in $\{x+\rho:\iota\in\mathbb{S}\wedge v(\rho)>v(x)\}$ if and only if $v(\varepsilon-\delta)>v(\delta-x\delta^{\prime})$.
We also recover (Corollary 3.1) the resonance-free part of Peran’s results [30, Corollary 2.4] on the field $\mathbb{T}_{\log}$ of logarithmic transseries [21]. In the resonant case of formal power series with exponents in an ordered field, we have a result (Theorem 3.3), and a counterexample (Section 3.3) to the convexity of the conjugacy problem in the non-resonant case (Corollary 2.1).
1 Groups of contracting derivations
Throughout the paper, we fix a field $C$ of characteristic 0, a non-empty directed set $(D,\leqslant)$ and a directed system $\mathcal{S}=(\Gamma_{d})_{d\in D}$, for the inclusion, of non-trivial ordered Abelian groups. We write $\Gamma$ for the direct limit of $\Gamma_{d}$.
Let $d\in D$. We have a field $\mathbb{S}_{d}:=C\left(\!(\Gamma_{d})\!\right)$ of Hahn series [22] with coefficients in $C$ and exponents in $\Gamma_{d}$. This is the ring, under pointwise sum and Cauchy product, of functions $f:\Gamma_{d}\longrightarrow C$ whose support $\operatorname{supp}f=\{g\in\Gamma_{d}:f(g)\neq 0\}$ is a well-ordered subset of $\Gamma_{d}$. It contains $C$ canonically, and we have canonical inclusions $\mathbb{S}_{d_{0}}\longrightarrow\mathbb{S}_{d_{1}}$ whenever $d_{0}\leqslant d_{1}$.
There is a notion of infinite sum for certain families in $\mathbb{S}_{d}$ called summable families (see [23, Section 3.1]), and the corresponding structure is a summability algebra in the sense of [9, Definition 1.27]. Let $\mathbb{S}$ be the direct limit of the directed system $(\mathbb{S}_{d})_{d\in D}$. This is a summability algebra for the direct limit summability structure, where a family is summable if and only if it takes values in an $\mathbb{S}_{d_{0}}$, for a $d_{0}\in D$, in which it is summable. A linear map $\mathbb{S}\longrightarrow\mathbb{S}$ (resp. $\mathbb{S}_{d}\longrightarrow\mathbb{S}_{d}$) that commutes with infinite sums is said strongly linear, and we write $\operatorname{Lin}^{+}(\mathbb{S})$ (resp. $\operatorname{Lin}^{+}(\mathbb{S}_{d})$) for the algebra under pointwise sum and composition of strongly linear maps on $\mathbb{S}$ (resp. $\mathbb{S}_{d}$).
Definition 1.1
A linear map $\phi:\mathbb{S}\longrightarrow\mathbb{S}$ is said regular if for all $d\in D$ we have $\phi(\mathbb{S}_{d})\subseteq\mathbb{S}_{d}$ and $\phi\upharpoonleft\mathbb{S}_{d}$ is strongly linear. We write $\operatorname{Lin}^{\mathcal{S}}(\mathbb{S})$ for the set of regular linear maps $\mathbb{S}\longrightarrow\mathbb{S}$.
[…]
A linear map $\phi:\mathbb{S}\longrightarrow\mathbb{S}$ is said contracting if $\phi(f)\prec f$ for all $f\in\mathbb{S}^{\times}$. Given $d_{0}\in D$, we write $\operatorname{Lin}_{\prec}^{+}(\mathbb{S}_{d_{0}})$ and $\operatorname{Lin}_{\prec}^{\mathcal{S}}(\mathbb{S})$ for the set of contracting strongly linear maps $\mathbb{S}_{d_{0}}\longrightarrow\mathbb{S}_{d_{0}}$ and the set of contracting regular maps $\mathbb{S}\longrightarrow\mathbb{S}$ respectively.
Lastly, we write $1$-$\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ for the group, under composition, of regular algebra automorphisms $\sigma$ of $\mathbb{S}$ such that $\sigma(f)\sim f$ for all $f\in\mathbb{S}$. Since each $C\operatorname{Id}_{\mathbb{S}_{d_{0}}}+\operatorname{Lin}^{+}_{\prec}(\mathbb{S}_{d_{0}})$ has evaluations in the sense of [9, Definition 2.3], and in view of the definition of regularity, so has $C\operatorname{Id}_{\mathbb{S}}+\operatorname{Lin}^{\mathcal{S}}_{\prec}(\mathbb{S})$. So [9, Theorem 2.14] applies and yields:
Proposition 1.1
The set $\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S})$ of contracting regular derivations on $\mathbb{S}$ is a group for the Baker-Campbell-Hausdorff operation
$\partial\ast\mathrm{d}:=\partial+\mathrm{d}+\frac{1}{2}\llbracket\partial,\mathrm{d}\rrbracket+\frac{1}{12}(\llbracket\partial,\llbracket\partial,\mathrm{d}\rrbracket\rrbracket-\llbracket\mathrm{d},\llbracket\partial,\mathrm{d}\rrbracket\rrbracket)+\cdots.$ (1)
Furthermore we have a group isomorphism
$\exp:(\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S}),\ast)$ $\longrightarrow$ $(1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S}),\circ)$
$\partial$ $\longmapsto$ $\operatorname{Id}_{\mathbb{S}}+\partial+\frac{1}{2}\partial\circ\partial+\frac{1}{6}\partial\circ\partial\circ\partial+\cdots.$
Hidden terms in (1) are $\mathbb{Q}$-linear combinations of iterated Lie brackets of lengths >3>3.
Remark 1.1
We will freely use the fact that if a family $(f_{i})_{i\in I}\in\mathbb{S}^{I}$ is summable and $f_{i}\preccurlyeq g$ for all $i\in I$, then $\sum_{i\in I}f_{i}\preccurlyeq g$. This follows from the fact that $\operatorname{supp}\sum_{i\in I}f_{i}\subseteq\bigcup_{i\in I}\operatorname{supp}f_{i}$, see [23, Section 3].
1.1 Contractive hull of a derivation
We now introduce a slight generalisation of the class of groups defined in [6, Section 7.2]. Let $\partial:\mathbb{S}\longrightarrow\mathbb{S}$ be a fixed regular derivation with kernel $\operatorname{Ker}(\partial)=C$. The contractive hull of $\partial$ is defined as the following subset of $\mathbb{S}$:
$\operatorname{Cont}(\partial):=\left\{f\in\mathbb{S}:f\partial\text{ is contracting}\right\}.$
Identifying each $f\in\operatorname{Cont}(\partial)$ with the regular contracting derivation $f\partial$, we obtain a Lie bracket $\llbracket\cdot,\cdot\rrbracket:\operatorname{Cont}(\partial)\times\operatorname{Cont}(\partial)\longrightarrow\operatorname{Cont}(\partial);(f,g)\mapsto f\partial(g)-\partial(f)g$ on $\operatorname{Cont}(\partial)$. It is easy to see that $\operatorname{Cont}(\partial)\partial$ is closed under sums of summable families. Thus $\operatorname{Cont}(\partial)$ is a group for the operation
$f\ast g:=f+g+\frac{1}{2}\llbracket f,g\rrbracket+\frac{1}{12}(\llbracket f,\llbracket f,g\rrbracket\rrbracket-\llbracket g,\llbracket f,g\rrbracket\rrbracket)+\cdots.$ (2)
The inverse of an $f\in\operatorname{Cont}(\partial)$ for $\ast$ is simply f-f. We also have the following consequences of [6, Lemmas 7.14 and 7.15 and Remark 7.19]. We give the proofs here for completion.
Lemma 1.1
[…]
Proof.
We may switch $f$ and $g$, so it suffices to show that $\llbracket f,g\rrbracket\prec f$. Since $g\partial$ is contracting, we have $g\partial(g)\prec g$, which means that $\partial(g)\prec 1$. We deduce that $f\partial(g)\prec f$. We also have $\partial(f)g=(g\partial)(f)\prec f$ since $g\partial$ is contracting. We deduce that $\llbracket f,g\rrbracket\prec f$. ∎
Lemma 1.2
For $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ with $f\neq g$, we have $f-g\succ\llbracket f,g\rrbracket$.
Proof.
We have $\llbracket f,g\rrbracket=\llbracket f-g,g\rrbracket$, so this follows from Lemma 1.1. ∎
In view of (2), Lemmas 1.1 and 1.2 and Remark 1.1, we obtain:
Corollary 1.1
[…]
Lemma 1.3
For all $f,g\in\operatorname{Cont}(\partial)$, we have
$f\ast g\ast(-f)$ $\sim$ $g\text{\quad and}$ (3)
$f\ast g\ast(-f)-g$ $\sim$ $\llbracket f,g\rrbracket.$ (4)
Proof.
If $f=0$ or $g=0$, then $f\ast g\ast(-f)=g$. Suppose that $f,g\neq 0$. If $\llbracket f,g\rrbracket=0$, then $f\ast g=f+g=g+f=g\ast f$ so $f\ast g\ast(-f)=g$. Suppose that $\llbracket f,g\rrbracket\neq 0$ and set $A:=\llbracket f,g\rrbracket$. So $A\prec g$ by Lemma 1.1. We see with Lemma 1.1 that $f\ast g=f+g+\frac{1}{2}A+\varepsilon$ for an $\varepsilon\prec A$. So
$f\ast g\ast(-f)$ $=$ $\left(f+g+\frac{1}{2}A+\varepsilon\right)\ast(-f)$
$=$ $g+\frac{1}{2}A+\varepsilon+\frac{1}{2}\left(\llbracket f,-f\rrbracket+\llbracket g,-f\rrbracket+\left\llbracket\frac{1}{2}A+\varepsilon,-f\right\rrbracket\right)+\cdots$
$=$ $g+\frac{1}{2}A+\varepsilon+\frac{1}{2}\left(\llbracket g,-f\rrbracket+\left\llbracket\frac{1}{2}A+\varepsilon,-f\right\rrbracket\right)+\cdots$
$=$ $g+A+B$
for $B=\varepsilon+\frac{1}{2}\left\llbracket\frac{1}{2}A+\varepsilon,-f\right\rrbracket+\cdots\prec A$ by Lemma 1.1. This shows that $f\ast g\ast(-f)\sim g$ and that $f\ast g\ast(-f)-g\sim A=\llbracket f,g\rrbracket$. ∎
Example 1.1
If $D=\{\bullet\}$, $\Gamma_{\bullet}=\Gamma=(\mathbb{Z},+,0,<)$, and $\partial=\frac{\mathrm{d}}{\mathrm{d}t}$ is the derivation with respect to $t$ on $C(\!(t)\!)=C(\!(\mathbb{Z})\!)$, then $\operatorname{Cont}(\partial)=\left\{f\in C(\!(\mathbb{Z})\!):f\prec t^{2}\right\}$.
1.2 Integration and asymptotic integration
From now on, given $f\in\mathbb{S}$ and $g\in\mathbb{S}^{\times}$, we sometimes write $f^{\prime}:=\partial(f)$ and $g^{{\dagger}}:=\partial(g)/g$. We make the assumption that $(\mathbb{S},\preccurlyeq,\partial)$ is an H-asymptotic field in the sense of [4, p 324]. In other words, we assume that for all $f,g\in\mathbb{S}^{\times}$ with $f,g\prec 1$, we have
[…]
Since $\operatorname{Ker}(\partial)+\mathbb{S}^{\prec}$ is the valuation ring of $(\mathbb{S},v)$, this means that $v$ is a differential valuation on $(\mathbb{S},\partial)$ in the sense of [34, Definition, p 4]. We then have well-defined maps
[…]
and the structure $\left(\Gamma,+,0,<,\mathord{\>{}^{{\dagger}}}\right)$ is called the asymptotic couple of $(\mathbb{S},\preccurlyeq,\partial)$ (see [4, p 325]). By [34, Theorem 4], it is an asymptotic couple in the sense of [4, p 273]. We have:
Lemma 1.4
[4, Lemma 6.5.4] The map $\mathord{{}^{\prime}}:\Gamma\setminus\{0\}\longrightarrow\Gamma$ is strictly increasing.
Given $f\in\mathbb{S}$, an asymptotic integral of $f$ in $(\mathbb{S},\preccurlyeq,\partial)$ is an element $A\in\mathbb{S}$ with $A^{\prime}\sim f$. Such an element may not exist. If such an element always exist, then $(\mathbb{S},\preccurlyeq,\partial)$ is said to be closed under asymptotic integration. In the case when $D$ is a singleton, being closed under asymptotic integration is equivalent [3, Lemma 1.7] to the existence of a strongly linear right inverse for $\partial$.
A crucial property of $(\mathbb{S},\preccurlyeq,\partial)$ is that [4, Theorem 9.2.1] there is at most one $\gamma\in\Gamma$ such that $\gamma\not\in(\Gamma\setminus\{0\})^{\prime}$. If such an element $\gamma$ exist, then we call it the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$. If no such $\gamma$ exists, then $\mathbb{S}$ is closed under asymptotic integration. Indeed, since $\operatorname{Ker}(\partial)=C$, for any $f\in\mathbb{S}\setminus\{0\}$ and any $g$ with $v(g^{\prime})=v(f)$, the element $cg$ is an asymptotic integral of $f$ where $c$ is the leading coefficient of $f/g^{\prime}$.
An important subset of $\Gamma$ is the psi-set
$\Psi:=\{\gamma^{{\dagger}}:\gamma\in\Gamma\wedge\gamma>0\}\subseteq\Gamma.$
Indeed, the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$ is either the maximum of $\Psi$ if this maximum exists, or the unique $\gamma\in\Gamma$ with $\Psi<\gamma<\{\gamma^{\prime}:\gamma\in\Gamma\wedge\gamma>0\}$ if $\Psi$ has no maximum ([4, Theorem 9.2.1 and Corollary 9.2.4]) and such an element exists. We will require the following technical lemma:
Lemma 1.5
Let $d_{0},d\in D$ with $d_{0}\leqslant d$ and let $\gamma\in\Gamma_{d}$ such that $\gamma^{\prime}$ is the pseudo-gap of $(\mathbb{S}_{d_{0}},\preccurlyeq,\partial)$. Then we have $\mathbb{N}\gamma+\Gamma_{d_{0}}\subseteq(\mathbb{N}\gamma+\Gamma_{d_{0}})^{\prime}$.
Proof.
Let $\alpha\in\Gamma_{d_{0}}$ and $n\in\mathbb{N}$. If $\alpha\notin(\Gamma_{d_{0}})^{\prime}$, then $\alpha=\gamma^{\prime}$. Since $n+1\neq 0$, we have $n\gamma+\alpha=((n+1)\gamma)^{\prime}\in(\mathbb{N}\gamma)^{\prime}$. Suppose now that $\alpha=\beta^{\prime}$ for a $\beta\in\Gamma_{d_{0}}$. If $n=0$ then $n\gamma+\alpha=\beta^{\prime}\in\Gamma_{d_{0}}^{\prime}$. We now assume that $n>0$. We have $\gamma^{\prime}\geqslant(\Gamma_{d_{0}}^{\neq})^{{\dagger}}$ and $\gamma<0$, so $\gamma^{{\dagger}}=\gamma^{\prime}-\gamma>(\Gamma_{d_{0}}^{\neq})^{{\dagger}}$. In particular $\gamma^{{\dagger}}>\beta^{{\dagger}}$, whence $(n\gamma)^{{\dagger}}=\gamma^{{\dagger}}>\beta^{{\dagger}}$. Thus $(n\gamma)^{\prime}+\beta>n\gamma+\beta^{\prime}=n\gamma+\alpha$, which means in view of the Leibniz rule that $(n\gamma+\beta)^{\prime}=n\gamma+\alpha$, so $n\gamma+\alpha\in\Gamma_{d}^{\prime}$.∎
Lemma 1.6
For $f\in\mathbb{S}$, we have $f\in\operatorname{Cont}(\partial)$ if and only if $v(f)+\Psi>0$.
Proof.
We have $f\in\operatorname{Cont}(\partial)$ if and only if $fg^{\prime}\prec g$ for all $g\in\mathbb{S}$, i.e. if and only if $f\prec\frac{1}{g^{{\dagger}}}$ for all $g\in\mathbb{S}\setminus C$. Since $(g^{-1})^{{\dagger}}=-g^{{\dagger}}\asymp g^{{\dagger}}$ for all $g\in\mathbb{S}$ and since $(c+\varepsilon)^{{\dagger}}\asymp\varepsilon^{\prime}\prec\varepsilon^{{\dagger}}$ for all $c\in C^{\times}$ and $\varepsilon\in\mathbb{S}$ with $\varepsilon\prec 1$, it is equivalent that $fg^{{\dagger}}\prec 1$ for all $g\in\mathbb{S}^{\prec}$, hence the result. ∎
Remark 1.2
Given a field extension $L/C$, there is a natural “strong extension of scalars” $\mathbb{S}\otimes^{+}_{C}L$ given as the direct limit of the system of fields $\left(L\left(\!(\Gamma_{d})\!\right)\right)_{d\in D}$. The map $\partial$ extends uniquely into a regular derivation $\partial_{L}:\mathbb{S}\otimes^{+}_{C}L\longrightarrow\mathbb{S}\otimes^{+}_{C}L$, and this extension preserves all relevant properties of $(\mathbb{S},\preccurlyeq,\partial)$. Namely $(\mathbb{S}\otimes^{+}_{C}L,\preccurlyeq,\partial_{L})$ is an H-asymptotic field with the same asymptotic couple as $(\mathbb{S},\preccurlyeq,\partial)$. So our results apply without change to $\mathbb{S}\otimes^{+}_{C}L$.
For $d\in D$, we write $\mathfrak{M}_{d}$ for the subset of series in $\mathbb{S}_{d}$ whose support is a singleton $\{\gamma\},\gamma\in\Gamma$ and whose value at $\gamma$ is $1$. So $\mathfrak{M}_{d}$ is a subgroup of $\mathbb{S}_{d}^{\times}$ and $v:(\mathfrak{M}_{d},\cdot,1,\succ)\longrightarrow(\Gamma_{d},+,0,<)$ is an isomorphism. Elements in $\mathfrak{M}=\bigcup_{d\in D}\mathfrak{M}_{d}$ are called monomials, and elements in $C\mathfrak{M}\subseteq\mathbb{S}$ are called terms. Given $f\in\mathbb{S}^{\times}$, there is a unique term $\operatorname{lead}(f)$ called the leading term of $f$ such that $f\sim\operatorname{lead}(f)$.
Remark 1.3
If $f,g\in\mathbb{S}$ and $(f_{i})_{i\in I}$ is a summable family in $\mathbb{S}$, then $\operatorname{supp}fg$ is contained in $\operatorname{supp}f+\operatorname{supp}g$, and $\operatorname{supp}\sum\limits_{i\in I}f_{i}$ is contained in $\bigcup\limits_{i\in I}\operatorname{supp}f_{i}$ (see [23, Section 3]). It follows that for $f,g\in\operatorname{Cont}(\partial)$ and for any submonoid $M\subseteq\Gamma$ with $\operatorname{supp}f\cup\operatorname{supp}g\subseteq M$ and $\partial(C(\!(M)\!))\subseteq C(\!(M)\!)$, the support of $f\ast g$ is contained in $M$.
We now define our notion of compatibility between asymptotic integration and the directed system. We will see that it is satisfied in most natural fields of generalised power series, such as transseries.
Definition 1.2
We say that asymptotic integration is regular on $\mathbb{S}$ if for all $d_{0}\in D$, there is a $d\in D$ with $d\geqslant d_{0}$ such that for each $f\in\mathbb{S}_{d_{0}}$ whose valuation is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$, there is a term $\tau\in C\mathfrak{M}_{d}$ with $\tau^{\prime}\sim f$ and $\tau^{\prime}\in\mathbb{S}_{d_{0}}$. We then say that such a $d$ is associated to $d_{0}$. We say that $\mathbb{S}$ has regular asymptotic integration if it is closed under asymptotic integration and asymptotic integration is regular on $\mathbb{S}$.
2 Conjugacy of derivations
We first consider the following approximation of the conjugacy problem: given $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ with $f\neq g$, when is there a $\varphi\in\operatorname{Cont}(\partial)$ such that $\varphi\ast g\ast(-\varphi)-f\prec f-g$? We say that a $\varphi\in\operatorname{Cont}(\partial)$ satisfying this is an asymptotic conjugating element for $(f,g)$, and we say that $f$ and $g$ are asymptotically conjugate if such an element exists.
Lemma 2.1
Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ with $f\neq g$. Then $f$ and $g$ are asymptotically conjugate if and only if $f\sim g$ and $\frac{f-g}{g^{2}}$ has an asymptotic integral $A$ in $\mathbb{S}$ with $v(A)+v(g)+\Psi>0$, their asymptotic conjugating elements are exactly the series $gA$ for such $A$.
Proof.
Let $d\in D$ with $f,g\in\mathbb{S}_{d}$. In view of Equation 3, a first necessary condition is that $f\sim g$. Suppose that $f\sim g$ and set $\delta:=f-g$. We want to find a $y\in\operatorname{Cont}(\partial)$ such that $y\ast g\ast(-y)-g-\delta\prec\delta$. Let $y\in\operatorname{Cont}(\partial)$. By Equation 4, we have $y\ast g\ast(-y)-g\sim\llbracket y,g\rrbracket$. Thus $y$ is an asymptotic conjugating element for $(f,g)$ if and only if
$y^{\prime}-g^{{\dagger}}y\sim\frac{-\delta}{g}.$
The solutions are of the form $gA$ where $A^{\prime}\sim\frac{-\delta}{g^{2}}$. Thus $f$ and $g$ are asymptotically conjugate if and only if $\delta/g^{2}$ has an asymptotic integral $A$ in $\mathbb{S}$ such that $gA$ lies in $\operatorname{Cont}(\partial)$. We conclude with Lemma 1.6. ∎
Lemma 2.2
Suppose that asymptotic integration is regular on $\mathbb{S}$. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ be asymptotically conjugate. There are a $d\in D$, an ordinal $\lambda>0$ and a strictly $\prec$-decreasing sequence $(\tau_{\gamma})_{\gamma<\lambda}$ of terms in $\operatorname{Cont}(\partial)\cap\mathbb{S}_{d}$ such that writing $\varphi_{\eta}:=\sum_{\gamma<\eta}\tau_{\gamma}$ for all $\eta\leqslant\lambda$, the sequence $(\varphi_{\eta}\ast g\ast(-\varphi_{\eta})-f)_{\gamma\leqslant\lambda}$ is strictly $\prec$-decreasing, and one of the following occurs:
  1. a)
    $\varphi_{\lambda}\ast g\ast(-\varphi_{\lambda})=f$.
  2. b)
    $v\left(\frac{f-\varphi_{\lambda}\ast g\ast(-\varphi_{\lambda})}{g^{2}}\right)$ is the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$.
Proof.
Let $d_{0}\in D$ such that $f,g\in\mathbb{S}_{d_{0}}$, and let $d\geqslant d_{0}$ be as in the definition of regular asymptotic integration, with respect to $d_{0}$. If there is a $\mathfrak{u}\in\mathfrak{M}_{d}$ such that $v(\mathfrak{u}^{\prime})$ is the pseudo-gap of $(\mathbb{S}_{d_{0}},\preccurlyeq,\partial)$, and $\mathfrak{u}^{\prime}\in\mathbb{S}_{d_{0}}$, then we set $\mathfrak{m}:=\mathfrak{u}$. If not, we set $\mathfrak{m}:=1$. By induction on an ordinal $\alpha$, we construct a strictly $\prec$-decreasing sequence $(\tau_{\gamma})_{\gamma<\alpha}$ of terms in $\operatorname{Cont}(\partial)\cap(C(\!(\Gamma_{d_{0}}+\mathbb{N}v(\mathfrak{m}))\!))$ such that defining $\varphi_{\eta},\eta\leqslant\alpha$ as above, the sequence $(\varphi_{\eta}\ast g\ast(-\varphi_{\eta})-f)_{\eta\leqslant\alpha}$ is strictly $\prec$-decreasing and that for all $\gamma<\alpha$ and all $\varepsilon\prec\tau_{\gamma}$, we have $(\varphi_{\gamma+1}+\varepsilon)\ast g\ast(-\varphi_{\gamma+1}-\varepsilon)-f\prec\varphi_{\gamma}\ast g\ast(-\varphi_{\gamma})-f$. The $\varepsilon$ here is a stand-in for $\varphi_{\alpha}-\varphi_{\eta}$ for various values of $\eta<\alpha$.
Let $\alpha$ such that for all $\eta<\alpha$, the sequence $(\tau_{\gamma})_{\gamma<\eta}$ is defined and satisfies the conditions. We first make a few observations. Let $\gamma\leqslant\alpha$ and write $g_{\gamma}:=\varphi_{\gamma}\ast g\ast(-\varphi_{\gamma})$. We have $g\sim g_{\gamma}$ by (3), so $\frac{f-g_{\gamma}}{g^{2}}$ and $\frac{f-g_{\gamma}}{g_{\gamma}^{2}}$ have the same symptotic integrals. Recall that $(f-g_{\rho})_{\rho\leqslant\alpha}$ is strictly $\prec$-decreasing. So given $\rho_{1}\leqslant\rho_{2}\leqslant\alpha$, if $\frac{f-g_{\rho_{i}}}{g^{2}}$ has an symptotic integral $A_{i},i\in\{1,2\}$, then we have $v(A_{1})+v(g_{\rho_{1}})+\Psi>0\Longrightarrow v(A_{2})+v(g_{\rho_{2}})+\Psi>0$. Thus, by Lemma 2.1, the series $f$ and $g_{\gamma}$ are asymptotically conjugate if and only if $v(\frac{f-g_{\gamma}}{g^{2}})$ is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$.
Let us now give the inductive proof. If $\alpha$ is a limit, including the case when $\alpha=0$, then we need only check that we have $(\varphi_{\gamma+1}+\varepsilon)\ast g\ast(-\varphi_{\gamma+1}-\varepsilon)\prec g_{\alpha}-f$ for all $\gamma<\alpha$ and $\varepsilon\prec\tau_{\gamma}$. But this follows by induction, noting that $\varphi_{\eta}-\varphi_{\gamma+1}\prec\tau_{\gamma}$ for all $\gamma<\eta<\alpha$. Suppose that $\alpha=\eta+1$ is a successor ordinal. If $g_{\eta}=f$, then, setting $\lambda:=\alpha$, we are done. Suppose that $g_{\eta}\neq f$. If $\mu:=v\left(\frac{f-g_{\eta}}{g^{2}}\right)$ is the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$, then we set $\lambda:=\alpha$ and we are done.
Suppose now that $\mu$ is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$. Since $\mathfrak{m}^{\prime}\in\mathbb{S}_{d_{0}}$, Remark 1.3 applies, and the set $\operatorname{Cont}(\partial)\cap(C(\!(\Gamma_{d_{0}}+\mathbb{N}v(\mathfrak{m}))\!))$ is closed under $\ast$. We see with Lemma 1.5 that $v(\frac{f-g_{\eta}}{g^{2}})$ is not a pseudo-gap in $(\mathbb{S}_{d},\preccurlyeq,\partial)$, so by regularity of asymptotic integration, there is a term $\tau$ in $\mathbb{S}_{d}$ with $\tau^{\prime}\in\mathbb{S}_{d}$ and $\tau^{\prime}\sim-\frac{f-g_{\eta}}{g^{2}}$. Note that $\tau$ is unique, and in view of Lemma 1.5, we have $\tau\in C^{\times}\mathfrak{M}_{d_{0}}\mathfrak{m}^{\mathbb{N}}$. So we may extend the sequence by setting $\tau_{\eta}:=\operatorname{lead}(g)\tau$, which by Lemma 2.1 lies in $\operatorname{Cont}(\partial)$. We have $\tau^{\prime}\prec(\frac{\tau_{\gamma}}{\operatorname{lead}(g)})^{\prime}$ for all $\gamma<\eta$ since $(f-g_{\gamma})_{\gamma<\alpha}$ is strictly $\prec$-decreasing. We deduce with Lemma 1.4 that $\tau\prec\frac{\tau_{\gamma}}{\operatorname{lead}(g)}$, whence $\tau_{\eta}\prec\tau_{\gamma}$ for all $\gamma<\eta$.
Thus $\varphi_{\alpha}$ is well-defined, and it remains to show that the conditions related to that sequence are satisfied. By Lemma 2.1 with $(g_{\eta},f)$ in the role of $(f,g)$, we have
$s\ast\varphi_{\eta}\ast g\ast(-\varphi_{\eta}\ast(-s))-f\prec\varphi_{\eta}\ast g\ast(-\varphi_{\eta})-f$
for all $s\in\mathbb{S}$ with $s\sim\tau_{\eta}$. We note with Corollary 1.1 that $\tau_{\eta}+\varphi_{\eta}$ is of the form $s\ast\varphi_{\eta}$ for $s:=(\tau_{\eta}+\varphi_{\eta})\ast(-\varphi_{\eta})\sim\tau_{\eta}$, so we have
$(\varepsilon+\tau_{\eta}+\varphi_{\eta})\ast g\ast(-(\varphi_{\eta}+\tau_{\eta}+\varepsilon))-f\prec\varphi_{\eta}\ast g\ast(-\varphi_{\eta})-f$
for all $\varepsilon\in\mathbb{S}$ with $\varepsilon\prec\tau_{\eta}$, as claimed. The sequence $(\tau_{\gamma})_{\gamma<\alpha}$ is strictly $\prec$-decreasing, hence injective. Since $\mathbb{S}_{d}$ is a set, there is an ordinal $\lambda$ at which the process stops, i.e. one of the cases of the lemma occurs. ∎
Note that the recularity of asymptotic integration plays a crucial role in the previous proof in allowing us to remain in $\mathbb{S}_{d}$ within the induction, so that $\varphi_{\lambda}$ be well-defined.
Proposition 2.1
Suppose that asymptotic integration is regular on $\mathbb{S}$. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ with $f\sim g$, and assume that $\mu:=v\left(\frac{f-g}{g^{2}}\right)$ satisfies $\mu>(-v(g)-\Psi)^{\prime}$ and lies above any pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$. Then $f$ and $g$ are conjugate in $\operatorname{Cont}(\partial)$.
Proof.
[…]
Theorem 2.1
Suppose that $(\mathbb{S},\preccurlyeq,\partial)$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$. Then $f$ and $g$ are conjugate in $\operatorname{Cont}(\partial)$ if and only if $f\sim g$ and $\mu:=v\left(\frac{f-g}{g^{2}}\right)$ satisfies $\mu>(-v(g)-\Psi)^{\prime}$.
Proof.
If $f$ and $g$ are conjugate, then they are asymptotically conjugate, so by Lemma 2.1, there is an $\alpha\in\Gamma\setminus\{0\}$ with $\alpha^{\prime}=\mu$ and $\alpha+v(g)+\Psi>0$. We deduce by Lemma 1.4 that $\alpha^{\prime}>(-v(g)-\Psi)^{\prime}$. Conversely, suppose that $\mu>(-v(g)-\Psi)^{\prime}$ and that $f\sim g$. By asymptotic integration, we find an $\alpha\in\Gamma$ with $\alpha^{\prime}=\mu$, whence $\alpha+v(g)+\Psi>0$ by Lemma 1.4. Since $(\mathbb{S},\preccurlyeq,\partial)$ has no pseudo-gap, we conclude with Proposition 2.1. ∎
Corollary 2.1
[…]
We will see in Section 3.3 that the conclusion of Corollary 2.1 does not hold in the presence of pseudo-gaps. We conclude with a more standard formulation of the conjugacy problem for derivations.
Theorem 2.2
Suppose that $(\mathbb{S},\preccurlyeq,\partial)$ has regular asymptotic integration. For $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ such that $f\sim g$ and $v\left(\frac{f-g}{g^{2}}\right)>(-v(g)-\Psi)^{\prime}$, the derivations $f\partial$ and $g\partial$ are conjugate over $\operatorname{Aut}(\mathbb{S})$, i.e. there is a $\sigma=\exp(h\partial)\in\operatorname{Aut}(\mathbb{S})$ with $\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}=f\partial$.
Proof.
By Theorem 2.1, there is an $h\in\operatorname{Cont}(\partial)$ with $(-h)\ast g\ast h=f$. Set $\sigma:=\exp(h\partial)$. Then $\sigma\circ\exp(g\partial)\circ\sigma^{\operatorname{inv}}=\exp(f\partial)$. Now the conjugation by $\sigma$ is a strongly linear algebra automorphism of $\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S})$ (see [9, Proposition 1.28]), while $\log$ is given as a power series of its argument with constant coefficients. So $f\partial=\log(\sigma\circ\exp(g\partial)\circ\sigma^{\operatorname{inv}})=\sigma\circ\log(\exp(g\partial))\circ\sigma^{\operatorname{inv}}=\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}$. ∎
3 Conjugacy of formal series
We now state the results in Section 2 in terms of conjugacy of series under composition.
3.1 Transseries
Let $\mathbb{T}$ denote the ordered field of logarithmic-exponential transseries [16, 12] together with its standard derivation $\partial:f\mapsto f^{\prime}$ and let $\mathfrak{M}$ denote the set of transmonomials in $\mathbb{T}$, i.e. $\mathfrak{M}$ is a specific section of the valuation $v:\mathbb{T}\longrightarrow\Gamma\cup\{\infty\}$. We identify each value $\gamma\in v(\mathbb{T}^{\times})$ with the corresponding transmonomial $\mathfrak{m}\in\mathfrak{M}$ with $v(\mathfrak{m})=\gamma$. Recall that $\mathbb{T}$ is a direct limit of Hahn fields $\mathbb{T}_{m,n},m,n\in\mathbb{N}$ where $\mathbb{T}_{m,n}=\mathbb{R}\left(\!(\mathfrak{M}_{m,n})\!\right)$ is the field of transseries with exponential depth $\leqslant m$ and logarithmic depth $\leqslant n$ (see [16, 18]). The directed system $(\mathfrak{M}_{m,n})_{m,n\in\mathbb{N}}$ has inclusions $\mathfrak{M}_{m,n}\longrightarrow\mathfrak{M}_{m^{\prime},n^{\prime}}$ whenever $n^{\prime}\geqslant n$ and $m^{\prime}\geqslant m+(n^{\prime}-n)$ (see [12, (2.7)]). The derivation $\partial$ is regular [12, Section 3]. The identity series is denoted $x$, and we have $x^{\prime}=1$, so $x^{{\dagger}}=x^{-1}$. Note that contrary to [31, 32, 30, 29], we take $x$ to be positive infinite as opposed to positive infinitesimal. There is a bijective morphism $\log:(\mathbb{T}^{>},\cdot,1,<)\longrightarrow(\mathbb{T},+,0,<)$ where $\mathbb{T}^{>}=\{f\in\mathbb{T}:f>0\}$. Given $n\in\mathbb{N}$, we denote by $\log_{n}x$ the $n$-th iterate of $\log$ applied at $x$. We write $\mathbb{T}_{\log}:=\bigcup\limits_{n\in\mathbb{N}}\mathbb{T}_{0,n}$ for the field of logarithmic transseries [21].
[…]
Let $\mathbb{S}$ denote the subset of $\mathbb{T}$ of flat transseries, i.e. series $f\in\mathbb{T}$ with $\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}$ for all $\mathfrak{m}\in\operatorname{supp}f$. So $\mathbb{S}=\{f\in\mathbb{T}:\operatorname{supp}f\subseteq\mathfrak{M}_{\preceq\!\!\preceq x}\}$ where $\mathfrak{M}_{\preceq\!\!\preceq x}=\{\mathfrak{m}\in\mathfrak{M}:\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}\}$. Note that $\mathfrak{M}_{\preceq\!\!\preceq}$ is a subgroup of $\mathfrak{M}$, so $\mathbb{S}$ is a subfield of $\mathbb{T}$. We write $\mathbb{S}^{>\mathbb{R}}:=\mathbb{S}\cap\mathbb{T}^{>\mathbb{R}}$ and $\mathbb{S}^{>}:=\mathbb{S}\cap\mathbb{T}^{>}$.
Example 3.1
The transmonomials $(\log_{2}x)^{-5}$, $x^{2}$ and $\operatorname{e}^{\sqrt{\log x}}$ are flat, whereas $\operatorname{e}^{(\log x)^{2}}$ and $\operatorname{e}^{x}$ are not.
Lemma 3.1
$\mathbb{S}$ is a differential subfield of $\mathbb{T}$
Proof.
It suffices to show that $\partial(\mathfrak{M}_{\preceq\!\!\preceq x})\subseteq\mathbb{S}$. We prove by induction on $m\in\mathbb{N}$ with $\mathfrak{m}\in\bigcup\limits_{n\in\mathbb{N}}\mathbb{T}_{m,n}$ that $\mathfrak{m}\in\mathfrak{M}_{\preceq\!\!\preceq x}\Rightarrow\mathfrak{m}^{\prime}\in\mathbb{S}$. This is clear if $m=0$ since $\mathbb{T}_{\log}^{\prime}\subseteq\mathbb{T}_{\log}$. If $\mathfrak{m}\in\mathbb{T}_{m+1,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x}$ and the result holds at $m$, then we have $\operatorname{supp}\mathfrak{m}^{\prime}\subseteq\mathfrak{m}\operatorname{supp}\log(\mathfrak{m})^{\prime}$ where $\log(\mathfrak{m})\in\mathbb{T}_{m,n+1}$. We conclude by induction.∎
Lemma 3.2
$\mathbb{S}$ is closed under composition.
Proof.
Let $f\in\mathbb{S}$ with $f>\mathbb{R}$. It suffices to show that $\mathfrak{M}_{\preceq\!\!\preceq x}\circ f\subseteq\mathbb{S}$. We prove by induction on $m\in\mathbb{N}$ with $\mathfrak{m}\in\bigcup\limits_{n\in\mathbb{N}}\mathbb{T}_{m,n}$ that $\mathfrak{m}\in\mathfrak{M}_{\preceq\!\!\preceq x}\Rightarrow\mathfrak{m}\circ f\in\mathbb{S}$. The group $\mathfrak{M}_{\preceq\!\!\preceq x}$ is closed under taking real powers, whence by [12, 6.8] so is $\mathbb{S}$. Moreover $\mathbb{S}$ is closed under $\log$. So the result holds if $m=0$. Suppose that $\mathfrak{m}\in\mathbb{T}_{m+1,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x}$ and that the result holds at $m$. Considering $\mathfrak{m}^{-1}$ if necessary, we may assume that $\mathfrak{m}$ is infinite. We have
$\mathfrak{m}\circ f=\exp(\log(\mathfrak{m})\circ f).$
By construction of $\exp$ on $\mathbb{T}$ (see [12, 1.5–2.7]), each element $\mathfrak{n}$ in $\operatorname{supp}\mathfrak{m}\circ f$ has the form $\mathfrak{n}=\mathfrak{u}\mathfrak{n}_{1}\cdots\mathfrak{n}_{k}$ where $\mathfrak{u}$ is the largest monomial of $\exp(\log(\mathfrak{m})\circ f)$, $k\in\mathbb{N}$, and $\mathfrak{n}_{1},\dots,\mathfrak{n}_{k}$ are in $\operatorname{supp}\log(\mathfrak{m})\circ f$. As $\log\mathfrak{m}\in\mathbb{T}_{m,n+1}$, the induction hypothesis gives $\mathfrak{n}_{1}\cdots\mathfrak{n}_{k}\in\mathbb{S}$. Lastly, note that $\mathfrak{u}^{{\dagger}}\asymp(\mathfrak{m}\circ f)^{{\dagger}}$. The chain rule [12, Proposition 6.3] gives $(\mathfrak{m}\circ f)^{{\dagger}}=f^{\prime}\mathfrak{m}^{{\dagger}}\circ f$. But $\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}$ and the composotion by $f$ on the right is valuation preserving, so $(\mathfrak{m}\circ f)^{{\dagger}}\preccurlyeq f^{{\dagger}}\preccurlyeq x^{-1}$. So $\mathfrak{u}\in\mathbb{S}$, whence $\mathfrak{m}\circ f\in\mathbb{S}$. We conclude by induction. ∎
We have a directed system $\mathcal{S}:=(\mathfrak{M}_{m,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x})_{m,n\in\mathbb{N}}$ and we write $\mathbb{S}_{m,n}=\mathbb{R}(\!(\mathfrak{M}_{m,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x})\!)=\mathbb{S}\cap\mathbb{T}_{m,n}$ for all $m,n\in\mathbb{N}$. We now focus on the directed system $\mathcal{S}$, H-field $(\mathbb{S},\preccurlyeq,\partial)$, and the group $(\operatorname{Cont}(\partial),\ast,0)$.
Lemma 3.3
The direct limits $\mathbb{S}$, $\mathbb{T}$ and $\mathbb{T}_{\log}$ have regular asymptotic integration.
Proof.
We prove this for $\mathbb{S}$, the proofs in the other cases being identical. Let $(m,n)\in\mathbb{N}^{2}$ and $f\in\mathbb{S}_{m,n}$. The pseudo-gap of $\mathbb{S}_{m,n}$ is $\gamma_{n}:=v((\log_{n+1}x)^{\prime})$ where $\log_{n+1}x\in\mathbb{S}_{m,n+1}\setminus\mathbb{S}_{m,n}$. Thus if $v(f)\neq\gamma_{n}$, then $f$ has an asymptotic integral in $\mathbb{S}_{m,n}$. If $v(f)=\gamma_{n}$, then $c\log_{n+1}x$ is an asymptotic integral of $f$ in $\mathbb{S}_{m,n+1}$ with derivative in $\mathbb{S}_{m,n}$, where $c\in C$ is the leading coefficient of $f$. We also deduce that there is no pseudo-gap in $\mathbb{S}$, so the conditions of Definition 1.2 hold for $(d_{0},d)=((m,n),(m,n+1))$. ∎
Lemma 3.4
We have $\operatorname{Cont}(\partial)=\{f\in\mathbb{S}:f\prec x\}$.
Proof.
We have $\Psi\geqslant v(x^{-1})$ by definition of $\mathbb{S}$, whence $v(x^{-1})=\min\Psi$, hence the result by Lemma 1.6. ∎
Given $f\in\mathbb{S}^{>\mathbb{R}}$, the right composition with $f$ is the map $\mathbb{S}\longrightarrow\mathbb{S}\>;\>g\mapsto g\circ f$. We say that a map $\sigma:\mathbb{S}\longrightarrow\mathbb{S}$ satisfies a chain rule if there is an $h\in\mathbb{T}$ such that for all $g\in\mathbb{S}$, we have
$\partial(\sigma(g))=h\sigma(\partial(g))$ (5)
Right compositions satisfy chain rules [12, Proposition 6.3]. We have a converse relation:
Lemma 3.5
Let $\sigma\in 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ satisfy a chain rule. If $\sigma(\log s)-\log s\prec 1$ for all $s\in\mathbb{S}^{>}$, then $\sigma$ is the right composition with $\sigma(x)$.
Proof.
Consider the contracting map $\phi:=\sigma-\operatorname{Id}_{\mathbb{S}}$. Let $\mathfrak{m}\in\mathfrak{M}$ with $\mathfrak{m}\succ 1$ and let $\mathfrak{n}\in\operatorname{supp}\log\mathfrak{m}$. We have $1\prec\mathfrak{n}\prec\mathfrak{m}$ by construction of $\log$ on $\mathbb{T}$. It follows since $(\mathbb{T},\preccurlyeq,\partial)$ is an H-asymptotic field that $\mathfrak{n}^{{\dagger}}\asymp(\mathfrak{n}^{-1})^{{\dagger}}\preccurlyeq(\mathfrak{m}^{-1})^{{\dagger}}\asymp\mathfrak{m}^{{\dagger}}$. This entails that $\log(\mathfrak{M}_{\preceq\!\!\preceq x})\subseteq\mathfrak{M}_{\preceq\!\!\preceq x}$, so $\log\mathbb{S}^{>}\subseteq\mathbb{S}$. Consider an $s\in\mathbb{S}^{>\mathbb{R}}$. We have
$\partial(\log(\sigma(s)))-\partial(\sigma(\log s))$ $=$ $\frac{\partial(\sigma(s))}{\sigma(s)}-h\sigma(\partial(\log s))$
$=$ $h\left(\frac{\sigma(\partial(s))}{\sigma(s)}-\sigma\left(\frac{\partial(s)}{s}\right)\right)$
$=$ $0.$
So $c:=\log(\sigma(s))-\sigma(\log s)\in\operatorname{Ker}(\partial)=C$. Our hypothesis on $\sigma$ gives $c\prec 1$, whence $c=0$, i.e. $\log(\sigma(s))=\sigma(\log s)$. An easy induction on the directed system $(\mathbb{S}_{m,n})_{m,n\in\mathbb{N}}$ shows that the right composition with $\sigma(x)$ is the only strongly linear automorphism of $\mathbb{S}$ which commutes with $\log$ and sends $x$ to $\sigma(x)$. So it must coincide with $\sigma$. ∎
Lemma 3.6
If $g\in\operatorname{Cont}(\partial)$, then $\exp(g\partial)$ is a right composition. Moreover, the derivation $\partial$ is contracting on $\mathbb{S}$ and $\exp(\partial)$ coincides with the right composition with $x+1$.
Proof.
Note that $\sigma:=\exp(g\partial)$ commutes with $g\partial$. Thus
$\partial\circ\sigma=\frac{1}{g}(g\partial)\circ\sigma=\frac{1}{g}\sigma\circ(g\partial)=\frac{\sigma(g)}{g}\sigma\circ\partial.$
Therefore $\sigma$ satisfies a chain rule.
Let $s\in\mathbb{S}^{>}$. We have $\sigma(\log s)-\log s\sim g(\log s)^{\prime}$. Since $g\in\operatorname{Cont}(\partial)$, we have $g(\log s)^{\prime}\prec(\log s)^{\prime}=s^{{\dagger}}$. But $s\in\mathbb{S}^{\times}$, so $s^{{\dagger}}\preccurlyeq x^{-1}$, so $\sigma(\log s)-\log s\prec 1$. Applying the previous lemma, we see that $\sigma$ is a right composition.
We know that $\partial$ is contracting by Lemma 3.4. Moreover $\exp(\partial)$ commutes with $\partial$, so it satisfies a chain rule. Therefore $\exp(\partial)$ is the right composition with $\exp(\partial)(x)=x+1+\partial(1)+\partial(\partial(1))+\cdots=x+1$. ∎
Theorem 3.1
The set $\mathcal{P}:=\{x+\delta\in\mathbb{S}:\delta\prec x\}$ is a group under composition, and the map
$\mathcal{E}:\operatorname{Cont}(\partial)$ $\longrightarrow$ $\mathcal{P}$
$f$ $\longmapsto$ $\exp(f\partial)(x)$
is a group isomorphism between $(\operatorname{Cont}(\partial),\ast,0)$ and $(\mathcal{P},\check{\circ},x)$.
Proof.
That $\mathcal{E}$ ranges in $\mathcal{P}$ follows from the fact that $\exp(\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S}))\subseteq 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$. Each $\exp(f\partial)$ for $f\in\operatorname{Cont}(\partial)$ commutes with $f\partial$, so it satisfies a chain rule. By Lemma 3.5, this implies that $\mathcal{E}$ is injective. Let us show that it is a surjective morphism. Let $f\in\mathcal{P}$. Considering $f^{\operatorname{inv}}$ if necessary, we may assume that $f>x$. By [19, Theorem 4.1], there is a series $V\in\mathbb{T}^{>\mathbb{R}}$ with $V^{\prime}\succ x^{-1}$ and $V\circ f=V+1$. We have
$(V^{\operatorname{inv}})^{{\dagger}}=\frac{(V^{\operatorname{inv}})^{\prime}}{V^{\operatorname{inv}}}=\left(\frac{1}{V^{\prime}x}\right)\circ V^{\operatorname{inv}}.$
Now $1/(V^{\prime}x)\prec 1$ so $(V^{\operatorname{inv}})^{{\dagger}}\prec 1$ by [24, Proposition 5.10].
We claim that $V^{{\dagger}}\preccurlyeq x^{-1}$. Indeed, we have $f-x>x^{-n}$ for a certain $n>1$. Note that $x^{n+1}\circ(x+x^{-n})>x^{n+1}+1$. The ordered group $(\mathbb{T}^{>\mathbb{R}},\circ,x,<)$ is a growth order group with Archimedean centralisers as a consequence of [8, Theorem 4.7]. So by the axiom GOG2 of [8, Section 2.1], for all $\varphi,\psi\in\mathbb{T}^{>\mathbb{R}}$ such that $\psi$ lies above all iterates of $x+x^{-n}$ and that $\varphi>\psi\circ\psi$, we have $\varphi\circ(x+x^{-n})\circ\varphi^{\operatorname{inv}}>\psi\circ(x+x^{-n})\circ\psi^{\operatorname{inv}}$. Here we apply this to $\psi=x^{n+1}$, and see that we must have $V\leqslant x^{2(n+1)}$. Thus $V^{{\dagger}}\preccurlyeq x^{-1}$. Now writing $V=V_{0}+V_{1}$ where $V_{0}\in\mathbb{S}$ and $V_{1}^{{\dagger}}\succ x^{-1}$, we have $V_{0}\circ f+V_{1}\circ f=V_{0}+1+V_{1}+1$ where $V_{0}\circ f-V_{0}-1\in\mathbb{S}$, so we must have $V_{1}\circ f-V_{1}=0$, and thus we may assume that $V=V_{0}\in\mathbb{S}$. By Lemma 3.6, we have
$f=(V^{\operatorname{inv}}\circ(x+1))\circ V=\sum_{i\in\mathbb{N}}\frac{(V^{\operatorname{inv}})^{(n)}\circ V}{i!}.$
We have $\frac{1}{V^{\prime}}\prec x$, so the regular derivation $\mathrm{d}:=\frac{1}{V^{\prime}}\partial$ on $\mathbb{S}$ is contracting by Lemma 3.4. An easy induction using the chain rule [12, Proposition 6.3] shows that $(V^{\operatorname{inv}})^{(i)}\circ V$ is the value of the $i$-th iterate of $\mathrm{d}$ at $x$, for all $i>0$. So $\exp(\mathrm{d})(x)=f$. We see with Lemma 3.6 that $\exp(\mathrm{d})$ is the right composition with $f$. This shows that $\mathcal{E}$ is surjective. Now $\exp:\operatorname{Der}^{\mathcal{S}}(\mathbb{S})\longrightarrow 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ is a group morphism, and $\sigma\mapsto\sigma(x)$ is a group morphism $\left(1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S}),\circ,\operatorname{Id}_{\mathbb{S}}\right)\longrightarrow(\mathcal{P},\check{\circ},x)$, so the result follows. ∎
Lemma 3.7
For all $\delta,\varepsilon\in\mathbb{T}^{\times}$ with $\delta,\varepsilon\prec x$ and $\delta\sim\varepsilon$ we have $1-x\delta^{{\dagger}}\sim 1-x\varepsilon^{{\dagger}}$.
Proof.
Write $\delta=\varepsilon+\iota$ where $\iota\prec\varepsilon$. So $1-x\delta^{{\dagger}}=1-x\varepsilon^{{\dagger}}-x\left(1+\iota/\varepsilon\right)^{{\dagger}}$ Recall that $\left(\Gamma,+,0,<,\mathord{\>{}^{{\dagger}}}\right)$ is an asymptotic couple, so we have $\left(1+\iota/\varepsilon\right)^{{\dagger}}\sim\left(\iota/\varepsilon\right)^{\prime}\prec\varepsilon^{{\dagger}}$ by [4, axiom AC3, p 273], hence the result. ∎
Theorem 3.2
Two series $x+\delta,x+\varepsilon\in\mathcal{P}\setminus\{x\}$ are conjugate in $\mathcal{P}$ if and only if $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$.
Proof.
Note that $\exp(f\partial)(x)-x\sim f$ for all $f\in\operatorname{Cont}(\partial)$. Recall that $v(x^{-1})=\min\Psi$. By Lemma 3.3, the direct limit $\mathbb{S}$ has regular asymptotic integration, so Theorem 2.1 applies. Now by Theorem 3.1 we can translate the conditions of conjugacy in Theorem 2.1, as the conjunction $\varepsilon\sim\delta$ and $v\left(\frac{\varepsilon-\delta}{\delta^{2}}\right)>v\left(\left(\frac{x}{\delta}\right)^{\prime}\right)$. Since $1-\delta x^{-1}\preccurlyeq 1$, the inequality $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$ entails that $\varepsilon\sim\delta$. We conclude with Lemma 3.7. ∎
Remark 3.1
For $\delta\in\mathbb{S}$ with $\delta\prec x$, since $\delta^{{\dagger}}\preccurlyeq x^{-1}$, we have $1-x\delta^{{\dagger}}\preccurlyeq 1$. Furthermore, we have $1-x\delta^{{\dagger}}\prec 1$ if and only if $\delta^{{\dagger}}\sim x^{-1}$, i.e. if and only if $\delta=xh$ for an $h\in\mathbb{S}$ with $h^{{\dagger}}\prec x^{-1}$. In all other cases, the series $x+\delta$ and $x+\varepsilon$ are conjugate if and only if $\varepsilon\sim\delta$.
Definition 3.1
Let $\mathcal{U}=(\Lambda_{d})_{d\in D}$ be a direct system of non-trivial ordered Abelian groups such that each $\Lambda_{d}$ for $d\in D$ is a subgroup of $\Gamma_{d}$ and that the morphisms $\Lambda_{d_{0}}\longrightarrow\Lambda_{d_{1}}$ for $d_{0}\leqslant d_{1}$ are restrictions of the morphisms $\Gamma_{d_{0}}\longrightarrow\Gamma_{d_{1}}$. Set $\mathbb{U}:=\underset{\longrightarrow}{\lim}_{d\in D}C\left(\!(\Lambda_{d})\!\right)$ and $\mathcal{P}_{\mathcal{U}}:=\{x+\delta:\delta\in\mathbb{U}\wedge\delta\prec x\}$. We say that $\mathcal{U}$ is a subsystem of $\mathcal{S}$ if
  1. a)
    $\partial$ is $\mathcal{U}$-regular, and
  2. b)
    $\mathcal{P}_{\mathcal{U}}$ is closed under composition.
For instance, the direct system $(\mathfrak{M}_{0,n})_{n\in\mathbb{N}}$ corresponding to the field $\mathbb{T}_{\log}$ of logarithmic transseries, is a subsystem of $\mathcal{S}$.
Proposition 3.1
Let $\mathcal{U}\subseteq\mathcal{S}$ be a subsystem and let $\mathbb{U}\subseteq\mathbb{S}$ denote the corresponding direct limit. Then $\exp(\operatorname{Cont}(\partial)\cap\mathbb{U})(x)=\mathcal{P}_{\mathcal{U}}$.
Proof.
For $f\in\operatorname{Cont}(\partial)\cap\mathbb{U}$, the series $\exp(f\partial)(x)=x+f+\frac{1}{2}ff^{\prime}+\cdots$ lies in $\mathbb{U}$ by $\mathcal{U}$-regularity of $\partial$. Conversely, let $\delta\in\mathbb{U}$ with $\delta\prec x$, and let $h\in\mathbb{S}$ with $x+\delta=\exp(h\partial)(x)$. Assume for contradiction that $h\not\in\mathbb{U}$. Recall that $\operatorname{supp}h$ is a well-ordered subset of $\Gamma_{d}$ for a $d\in D$. So there is a least element $\gamma_{0}\in\operatorname{supp}h\setminus\Lambda_{d}$. Write $\mathfrak{m}_{0}$ for the corresponding element of $\mathfrak{M}_{d}$. Let $\varphi$ denote the element of $\mathbb{S}_{d}$ with $\operatorname{supp}\varphi:=\{\gamma\in\operatorname{supp}h:\gamma<\gamma_{0}\}$. So $\varphi\in\mathbb{U}_{d}$ and $h=\varphi+\psi$ where $\psi\sim h(\gamma_{0})\mathfrak{m}_{0}$. By Lemma 1.1, the series $\varepsilon:=h\ast(-\varphi)=\psi+\frac{1}{2}\llbracket\psi,-\varphi\rrbracket+\cdots$ satisfies $\varepsilon\sim h(\gamma_{0})\mathfrak{m}_{0}$, so $\exp(\varepsilon\partial)(x)-x\sim h(\gamma_{0})\mathfrak{m}_{0}$. In particular, we have $\exp(\varepsilon\partial)(x)\not\in\mathbb{U}$. But $\exp(\varepsilon\partial)(x)=\exp(-\varphi\partial)(x)\circ\exp(h\partial)(x)\in\mathcal{P}_{\mathcal{U}}$ by Definition 3.1(b): a contradiction. ∎
Corollary 3.1
For all $n\in\mathbb{N}$, two series $x+\delta,x+\varepsilon\in\mathcal{P}\cap\mathbb{T}_{0,n}\setminus\{x\}$ are conjugate in $\mathcal{P}\cap\mathbb{T}_{0,n+1}$ if and only if $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$. In particular, two series $x+\delta,x+\varepsilon\in\mathcal{P}\cap\mathbb{T}_{\log}\setminus\{x\}$ are conjugate in $\mathcal{P}\cap\mathbb{T}_{\log}$ if and only if $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$.
Proof.
Recall by Lemma 3.3 that $\mathbb{T}_{\log}$ has regular asymptotic integration. The pseudo-gap $v((\log_{n+1}x)^{\prime})$ of $\mathbb{T}_{0,n+1}$ is strictly below the set $v(\mathbb{T}_{0,n})$. So we may apply Proposition 2.1 in $\operatorname{Cont}(\partial\restriction\mathbb{T}_{0,n+1})$. Then, as in the proof of Theorem 3.2, we translate the inequalities using the isomorphism from Proposition 3.1. ∎
Example 3.2
Another example of subsystem of $\mathcal{S}$ is the direct system corresponding to flat grid-based transseries [16, 24, 18]. The valued differential field of flat grid-based transseries has regular asymptotic integration, so one obtains the same conditions for conjugacy of parabolic flat grid-based transseries.
Example 3.3
Consider the subsystem $\left(\frac{1}{n!}\mathbb{Z}\right)_{m,n\in\mathbb{N}}$ of $\mathcal{S}$. This corresponds to formal Puiseux series over $\mathbb{R}$. Here we have a pseudo-gap $v(x^{-1})$, so we only obtain the sufficient but non-necessary conditions as in Theorem 3.3.
Remark 3.2
Our results up to and excluding Theorem 3.1 apply when replacing $\mathbb{T}$ with the larger, class-sized field of $\omega$-series [5, Section 4.5]. This is an totally ordered direct limit of (set-sized) fiels of generalised series indexed by all ordinals. Although it has a pseudo-gap, asymptotic integration is regular with respect to this directed system.
3.2 Formal power series with exponents in a field
Let $C$ be an ordered field and let $\Gamma$ denote its underlying ordered additive group. Then the field $\mathbb{K}=C\left(\!(\Gamma)\!\right)$ is endowed with a standard derivation $\partial$ given by $\partial(f)(c)=(c+1)f(c+1)$ for all $(f,c)\in\mathbb{K}\times C$. We write each element $f$ of $\mathbb{K}$ as a formal series $f=\sum_{c\in C}f(c)x^{c}$, so $\partial(f)=\sum_{c\in C}cf(c)x^{c-1}$, and $\partial(x)=1$. Note that $\operatorname{Cont}(\partial)=\{f\in\mathbb{K}:f\prec x\}$ and that the H-asymptotic field $(\mathbb{K},\preccurlyeq,\partial)$ is grounded in the sense of [4, p 326], i.e. the set $\Psi$ has a maximum $v(x^{-1})$, which is thus the pseudo-gap of $(\mathbb{K},\preccurlyeq,\partial)$.
We showed [7, Proposition 6.6] that $\exp(\operatorname{Cont}(\partial))(x)$ is the group
$\mathcal{P}_{C}:=\{x+\delta:\delta\in\mathbb{K}\wedge\delta\prec x\}$
of parabolic series in $\mathbb{K}$ under the formal composition law of [7, Section 6.1]. In view of Proposition 2.1, the same arguments as in the proof of Theorem 3.2 entail that for all $\delta,\varepsilon\in\mathbb{K}$ with $\delta,\varepsilon\prec x$ and $\varepsilon\sim\delta$, the series $x+\delta$ and $x+\varepsilon$ are conjugate in $\mathcal{P}_{C}$ if $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$ and $v\left(\frac{\varepsilon-\delta}{\delta^{2}}\right)$ lies above the pseudo-gap of $\mathbb{K}$, i.e. if $v\left(\frac{\varepsilon-\delta}{\delta^{2}}\right)+v(x^{-1})>0$. This translates to $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}}),x\delta^{2}$. If $\delta\prec x^{-1}$, then as in Remark 3.1, we have $v(1-x\delta^{{\dagger}})=0$, so $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$ holds because $\varepsilon\sim\delta$. If $\delta\succcurlyeq x^{-1}$, then $x\delta^{2}\succcurlyeq\delta$ so $\varepsilon-\delta\prec x\delta^{2}$ holds because $\varepsilon\sim\delta$. Therefore:
Theorem 3.3
For all $\delta,\varepsilon\in\mathbb{K}$ with $\delta,\varepsilon\prec x$ with $\varepsilon\sim\delta$, the series $x+\delta$ and $x+\varepsilon$ are conjugate in $\mathcal{P}_{C}$ if $\delta\succcurlyeq x^{-1}$ and $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$ or $\delta\prec x^{-1}$ and $\varepsilon-\delta\prec x\delta^{2}$.
3.3 A case of resonance
We conclude by giving a simple example of resonance. Consider the field $\mathbb{K}$ above for a given ordered field $C$. For all $c<0$ in $C$ and $\delta\in\mathbb{K}$ with $v(\delta)=v(x^{c})$, we write $f_{\delta}:=x+1+\delta$, and set $f_{0}:=x+1$. The pseudo-gap of $\mathbb{K}$ is $v(x^{-1})$. Thus for $\delta\in\mathbb{K}$ with $v(\delta)=v(x^{-1})$, the series $f_{0}$ and $f_{\delta}$ are not asymptotically conjugate, whence not conjugate. However, for $c\in(-1,0)$ in $C$, and $\varepsilon:=x^{c}$, the series $f_{0}$ and $f_{\varepsilon}$ are asymptotically conjugate, and the approximative conjugacy method of Lemma 2.1 (translated via the exponential map $\mathcal{E}:\operatorname{Cont}(\partial)\longrightarrow\mathcal{P}_{C}$) gives an asymptotic conjugating element $\varphi:=x+\frac{1}{c+1}x^{c+1}$. Using formal Taylor expansions (see [7, Section 6.1]) of up to order $2$, one obtains
$\varphi\circ f_{\varepsilon}\circ\varphi^{\operatorname{inv}}=\varphi\circ\left(\varphi^{\operatorname{inv}}+1+x^{c}-\frac{c}{c+1}x^{2c}+\cdots\right)=f_{\iota}$
[…]
Acknowledgments
We thank Daniel Panazzolo and Jean-Philippe Rolin for their answers to our questions.